# AfroXLMR-Social: Adapting Pre-trained Language Models for African Languages Social Media Text

### v2

@@
Tadesse Destaw Belay1, Israel Abebe Azime2, Ibrahim Said Ahmad3,4, David Ifeoluwa Adelani5,6
Idris Abdulmumin7
, Abinew Ali Ayele8, Shamsuddeen Hassan Muhammad4,9,
Seid Muhie Yimam10
1Instituto Politécnico Nacional, 2Saarland University, 3Northeastern University, 4Bayero University Kano, 5Mila, McGill University,
6Canada CIFAR AI Chair, 7University of Pretoria, 8Bahir Dar University, 9Imperial College London, 10University of Hamburg,
Contact: tadesseit@gmail.com
AfroXLMR-Social: Adapting Pre-trained Language Models
for African Languages Social Media Text
Tadesse Destaw Belay1, Israel Abebe Azime2, Ibrahim Said Ahmad3,4, David Ifeoluwa Adelani5,6 Idris Abdulmumin7, Abinew Ali Ayele8, Shamsuddeen Hassan Muhammad4,9, Seid Muhie Yimam10 1Instituto Politécnico Nacional, 2Saarland University, 3Northeastern University, 4Bayero University Kano, 5Mila, McGill University, 6Canada CIFAR AI Chair, 7University of Pretoria, 8Bahir Dar University, 9Imperial College London, 10University of Hamburg, Contact: tadesseit@gmail.com
Language models built from various sources are the foundation of today’s NLP progress. However, for many low-resource languages, the diversity of domains is often limited—more biased to a religious domain, which impacts their performance when evaluated on distant and rapidly evolving domains such as social media. Domain adaptive pre-training (DAPT) and task-adaptive pre-training (TAPT) are popular techniques to reduce this bias through continual pre-training for BERT-based models, but they have not been explored for African multilingual encoders. In this paper, we explore DAPT and TAPT continual pertaining approaches for the African languages social media domain. We introduce AfriSocial—a large-scale social media and news domain corpus for continual pre-training on several African languages. Leveraging AfriSocial, we show that DAPT consistently improves performance on three subjective tasks: sentiment analysis, multi-label emotion, and hate speech classification, covering 19 languages from 1% to 30% F1 score. Similarly, leveraging TAPT on one task data improves performance on other related tasks. For example, training with unlabeled sentiment data (source) for a fine-grained emotion classification task (target) improves the baseline results by an F1 score ranging from 0.55% to 15.11%. Combining these two methods (i.e. DAPT + TAPT) further improves the overall performance.
@@ 1 Introduction
Pre-trained language models (PLMs) are initially trained on vast and diverse corpora, including encyclopedias and web content Conneau et al. (2020a); Chiang et al. (2022). Subsequently, these pre-trained models are used in supervised training for a specific Natural Language Processing (NLP) task by further finetuning. Fine-tuned PLMs achieve strong performance across many tasks and datasets from various sources Wu et al. (2024). However, this raises a question: Do PLMs function universally across domains, or does continual training of PLMs with domain-specific data offer better performance?
@@
Figure 1: Continual pretraining illustration. A general-purpose pretrained model, such as AfroXLMR, is first adapted to a social domain, resulting in AfroXLMR-Social. This model then undergoes Cross-task TAPT using sentiment analysis, emotion, and hate speech data without the labels for further fine-tuning.
@@
While some studies have shown the benefit of continual pretraining on a domain-specific unlabeled data Lee et al. (2019); Gururangan et al. (2020), one domain may not be generalizable to other domains and languages. Moreover, it is unknown how the benefit of continual pretraining may vary with factors like the amount of unlabeled corpus, the source domain itself, the evaluation task, the resource richness of the target languages, and the trained target model Gururangan et al. (2020). This raises the question of whether pretraining on a corpus more directly tied to the task can further improve performance. This work addresses these questions by continual domain adaptive pertaining (DAPT) and task adaptive pretraining (TAPT) on the downstream subjective NLP tasks in a low-resource language setup. We consider the social media domain X and News for a continual pretraining from a high-performing multilingual baseline model, AfroXLMR Alabi et al. (2022), a continual pretrained model for African languages from XLM-RoBERTa Conneau et al. (2020a). We explore subjective NLP task results from baseline (the base model result), DAPT, TAPT, and DAPT + TAPT, on a smaller but directly domain- and task-relevant unlabeled corpus. Our result shows that DAPT and TAPT highly benefit from similar source NLP tasks. Figure 1 shows the general high-level continual pre-training illustration. In summary, our contributions are:
@@
  • Presenting AfriSocial, a new quality domain-specific corpus for 14 African languages, collected from the social domain (X) and news.
@@
  • A through analysis of domain and task adaptive continual pretraining across subjective NLP tasks for low-resource languages.
@@
  • Achieving state-of-the-art results in the evaluated NLP tasks and publicly making AfriSocial social-domain corpus and AfroXLMR-Social pretrained models for developing low-resource languages.
@@ 2 Related Work
Domain in NLP Language models (LMs) pretrained on text from various sources are the foundation of today’s NLP. Domain adaptation in NLP refers to enhancing the performance of a model using similar domain data (target domain) by leveraging knowledge from an existing domain (source domain) Ramponi and Plank (2020). Domain refers to different implicit clusters of text representations in pretrained LMs, such as news articles, social media, medical texts, or legal documents Aharoni and Goldberg (2020); Shi et al. (2024). Each domain has its unique characteristics, vocabulary, and writing style, which can affect the performance of NLP models when applied to new or unseen domains. Therefore, a similar domain means the text source from which the pre-trained model was made, similar to the target NLP task data source.
@@
Continual Pre-training of LMs There are several techniques for the downstream NLP task improvements, such as general-purpose pre-training Wang et al. (2023), language-adaptive continual pre-training Alabi et al. (2022), domain-adaptive pre-training Gururangan et al. (2020), task-adaptive pre-training Alabi et al. (2022), and data augmentation Zhang et al. (2024). Prior works have shown the benefit of continual pretraining using a domain-specific corpus Gururangan et al. (2020) and training LMs in a specific domain from scratch Huang et al. (2020). However, continual pretraining is arguably more cost-effective since it is a continuous pretraining from the existing base language model.
@@
Domain Adaptive Pre-training (DAPT) Domain Adaptive pre-training (DAPT) is straightforward, continuing pretraining a model on a corpus of unlabeled domain-specific text Aharoni and Goldberg (2020). DAPT techniques handle discrepancies between the source (pre-training) and target (fine-tuning) domains. Traditional fine-tuning often yields suboptimal results when pre-trained models encounter data that diverges significantly from their training data. DAPT techniques reduce this mismatch by aligning data distributions, ensuring the model can generalize better in a novel setting. Lee et al. (2019) considers a single domain at a time and uses a language model pretrained on a smaller and less diverse corpus than the most varied and multilingual language models.
@@
Task Adaptive Pre-training (TAPT) Task-adaptive pretraining (TAPT) refers to pretraining on the unlabeled training set for a similar task-specific data Gu et al. (2024). Compared to DAPT, TAPT solely leverages the training data of the similar downstream task for continuous pre-training. It uses a far smaller pretraining corpus that is much more task-relevant (assuming that the training set represents aspects of the task well). This makes TAPT much less expensive to run than DAPT. There are also a combination of various techniques, such as DAPT followed by TAPT, which is beneficial for end-task performance Gururangan et al. (2020).
@@
Language Adaptive Pre-training (LAPT) LAPT is also called language-adaptive fine-tuning (LAFT) Alabi et al. (2022). It mainly focuses on adapting a pre-trained language model to a specific language(s) using any language-specific corpus. This is the most common approach: collecting any corpus for fine-tuning language models without considering the domain Yu and Joty (2021); Alabi et al. (2022). LAPT is a vital step to improve language understanding and representation, especially for low-resource languages Wang et al. (2023). However, in the primary studies that investigated LAPT with various language-specific corpora, the impacts of the text sources on specific NLP tasks are unexplored.
@@
Although XLMR Conneau et al. (2020a) pretraining corpus is derived from multiple sources and languages, it has not yet been explored whether these sources are diverse enough to generalize in a specific domain and task. This leads to asking whether subjective tasks related to social media text can be understood with this generic model. Towards this end, we explore further a continual training of DATP and TAPT from AfroXLMR Alabi et al. (2022) and evaluating the impacts on highly subjective NLP tasks such as sentiment analysis, multi-label emotion, and hate speech classification.
@@ 3 AfriSocial: Domain Adaptation Corpus
In computational linguistics, domain boundaries can be defined through various dimensions including content, style, and purpose Plank (2016). In this work, we define the social domain based on two key factors: 1) Convergent textual characteristics, X and news sources provide a public discourse, comments, reactions, and conversational linguistic patterns that facilitate social interactions, and cultural expressions and 2) Functional similarity in downstream tasks, the selected evaluation task datasets are sourced from the two sources, shown in Table 1. This grouping is empirically supported by Derczynski et al. (2016), which demonstrated comparable performance patterns when evaluating entity recognition models across X and news data, suggesting underlying linguistic commonalities despite surface differences.
@@
To be specific, a domain-specific corpus is vital for adapting pre-trained language models into a specific domain, such as a health-specific domain. Likewise, the social media corpus is limited even to a highly resourced language.
Dataset name Task name # lang. Data Sources # Classes
AfriSenti Muhammad et al. (2023); Yimam et al. (2020) Sentiment analysis 14 News, social media 3
AfriEmo Muhammad et al. (2025b); Belay et al. (2025a) Emotion analysis 17 News, social media 6
AfriHate Muhammad et al. (2025a) Hate speech detection 15 News, social media 3
AfriSocial (new domain specific corpus) Unlabeled 14 X and news -
Table 1: List of datasets. Social media sources include posts/comments from YouTube and X. The AfriSenti dataset class labels are positive, negative, and neutral. The AfriEmo dataset labels are six basic emotions (anger, disgust, fear, joy, sadness, and surprise) in a multi-label annotation. AfriHate labels are abuse, hate, and neutral.
@@ 3.1 Data Sources Selection
Based on our assessment, most of the NLP datasets of African languages are sourced and annotated from X and news domain, as sufficient text can be found in these two sources. As shown in Table 1, subjective NLP tasks for African languages, such as sentiment analysis, multi-label emotion, and hate speech classification datasets, are sourced from X and news. The AfriSocial corpus is sourced from similar domains to improve these subjective tasks further, X and News. There is an X domain corpus and model for high-resource languages such as XLM-T Barbieri et al. (2022) to evaluate and improve task datasets sourced from X. However, no available corpora specialize in the social domain for low-resource African languages. More details about the data collections are presented in Appendix I.
@@ 3.2 Pre-processing and Quality Control
We apply better LID tools for each language. For example, for the language mixing problem of the existing corpus mentioned in Sec §3, we used separate LID tools, GeezSwitch [^1] https://pypi.org/project/geezswitch/ to handle Ethiopic script languages (Amharic and Tigrinya) and pycld3222https://pypi.org/project/pycld3/ for the supported Latin and other script languages at the sentence level.
@@
The same approach as language identification, we used proper tools for each language to segment into sentences. For the Ethiopic script language, we used amseg tool Yimam et al. (2021), and for other Latin script languages, we used NLTK Bird and Loper (2004).
@@
We exclude short sentences, only URL lines, and anonymize personally identifiable information (PII) such as usernames starting with the @ symbol, and email addresses. We pay special attention to ensuring that the available evaluation task data (sentiment, emotion, and hate speech) do not appear in the AfriSocial corpus before and after processing. De-duplication is applied if a near-similar instance is present, excluding it from AfriSocial, not from the annotated dataset. Table 2 shows the AfriSocial corpus statistics with their sources and number of sentences.
4 Evaluation Tasks and Datasets
We select subjective NLP tasks for our evaluation based on the following reasons. 1) Subjective tasks face more disagreement during annotations, leading to less performance in the evaluation, especially for low-resource languages Fleisig et al. (2023). As we can see from the SemEval-2025 Task 11 Muhammad et al. (2025c), an emotion detection shared task covering 32 languages, low-resource languages are not well explored, and the lowest results are from African languages. 2) African language subjective tasks are sourced from X and news, as shown in Table 1. These sources align with the same domain as the AfriSocial corpus. The three subjective tasks for our evaluation are sentiment analysis, multi-label emotion, and hate speech classification. We keep the original train-test split of all evaluation datasets throughout our experiment for later comparison with benchmark results.
4.1 AfriSenti: Sentiment Analysis Dataset
AfriSenti Muhammad et al. (2023) is a sentiment analysis dataset across 14 African languages. The data is sourced from X (formerly Twitter) and annotated in one of the three sentiment classes: positive, negative, and neutral. From 14 languages, the two languages, Oromo (orm) and Tigrinya (tir) have only test sets.
4.2 AfriEmo: Multi-label Emotion Dataset
SemEval-2025-Task 11 Muhammad et al. (2025c) is an emotion dataset that covers 32 languages, from diverse domains such as social media platforms (X, Reddit, YouTube, and others) and news. The AfriSocial domain-specific corpus targets low-resource African languages; we target the African languages emotion dataset from the SemEval-2025 Task 11, which we call AfriEmo. It covers 17 African languages from the 32 languages. This dataset is annotated in a multi-label approach - a text might have any combination (none, one, some, or all) of emotion labels from a given set of emotions (anger, disgust, fear, joy, sadness, and surprise).
4.3 AfriHate: Hate Speech Dataset
AfriHate Muhammad et al. (2025a) is a multilingual hate and abusive speech dataset in 15 African languages sourced from X. Each text is categorized into one of the abusive, hate, or neutral labels. The languages covered in the corresponding evaluation datasets, such as language name, ISO code, countries/regions spoken, language family, and writing script, are presented in the Appendix (A, B).
5 Language Models
5.1 Encoder-only Language Models
Multilingual encoder-only pretrained language models (PLMs) such as XLM-R Conneau et al. (2020a) and mBERT Devlin et al. (2019) have shown impressive capability on many languages for a variety of downstream tasks. They are also often used to initialize checkpoints to adapt to other languages, such as AfroXLMR Alabi et al. (2022), which is initialized from XLM-R to specialize in African languages. We evaluate popular multilingual and African-centric PLMs such as AfroLM Dossou et al. (2022) and AfriBERTa Ogueji et al. (2021) and found that AfroXLMR is better for the targeted evaluation datasets as it covers more African languages. We make AfroXLMR our benchmarks and further train on the AfriSocial domain and the selected evaluation tasks. The AfroXLMR plus DAPT with the AfriSocial domain-specific corpus gives us AfroXLMR-Social. The detailed hyperparameters of the continual training are shown in Appendix D.
5.2 Large Language Models
We compare our DAPT and TAPT approach results from AfroXLMR with state-of-the-art open source LLMs such as Llama 3 Grattafiori et al. (2024), Gemma 2 Team et al. (2024), Mistral Jiang et al. (2024), and proprietary LLMs such as Gemini 1.5 Georgiev et al. (2024) and GPT-4o Hurst et al. (2024). For AfriSenti and AfriHate task results, we use the LLMs benchmark results from the AfroBench Ojo et al. (2025) leaderboard. For the AfriEmo task, we use LLM results from the SemEval-2025 Task 11 datasets papers Muhammad et al. (2025b); Belay et al. (2025a). The detailed versions of the LLMs are presented in the Appendix E.
6 Experiment Results
6.1 Domain Adaptive Pre-training (DAPT)
The domain-adaptive pretraining (DAPT) approach is straightforward; we continue pretraining from the strong baseline language model (AfroXLMR) using the domain-specific AfriSocial corpus in a multilingual setup. Baseline results from (AfroXLMR) and after applying DAPT results are presented in Table 3.
@@ 6.1 Domain Adaptive Pre-training (DAPT)
Lang. AfroXLMR DAPT Lang. AfroXLMR DAPT Lang. AfroXLMR DAPT
@@
Baseline: As a baseline, we use the results from AfroXLMR Alabi et al. (2022), which perform better than other BERT-based encode-only models across targeted datasets since it includes more African languages during pre-training.
Results: The results before and after DAPT are shown in Table 3, AfroXLMR and DAPT columns, respectively. We observe that DAPT improves over the baseline in all languages and datasets, demonstrating the benefit of DAPT when the target domain is relevant. It consistently improves over the baseline models for each language. It suggests that continual pretraining on a small, quality, domain-relevant dataset is important for subjective tasks from the same domain.
6.2 Task Adaptive Pretraining (TAPT)
Similar to DAPT, TAPT consists of a second phase of continual pertaining. TAPT is a cross-task transfer across datasets, which refers to finetuning on the unlabeled data of the non-targeted task during evaluation. We explore the TAPT approach by directly applying it to the base model and after DAPT. For example, if we make TAPT for the AfriSenti task, we further fine-tune the base model and DAPT model using the unlabeled data of the AfriEmo and AfriHate datasets separately. Compared to DAPT, the task-adaptive approach strikes a different trade-off: it uses a far smaller pretraining corpus, assuming the training set represents aspects of the target task. This makes TAPT much less expensive to run than DAPT.

### v3

@@
Abstract
Language models built from various sources are the foundation of today’s NLP progress. However, for many low-resource languages, the diversity of domains is often limited, more biased to a religious domain, which impacts their performance when evaluated on distant and rapidly evolving domains such as social media. Domain adaptive pre-training (DAPT) and task-adaptive pre-training (TAPT) are popular techniques to reduce this bias through continual pre-training for BERT-based models, but they have not been explored for African multilingual encoders. In this paper, we explore DAPT and TAPT continual pre-training approaches for African languages social media domain. We introduce AfriSocial, a large-scale social media and news domain corpus for continual pre-training on several African languages. Leveraging AfriSocial, we show that DAPT consistently improves performance (from 1% to 30% F1 score) on three subjective tasks: sentiment analysis, multi-label emotion, and hate speech classification, covering 19 languages. Similarly, leveraging TAPT on the data from one task enhances performance on other related tasks. For example, training with unlabeled sentiment data (source) for a fine-grained emotion classification task (target) improves the baseline results by an F1 score ranging from 0.55% to 15.11%. Combining these two methods (i.e. DAPT + TAPT) further improves the overall performance. The data and model resources are available at HuggingFace [^1] https://huggingface.co/tadesse/AfroXLMR-Social.
@@ Abstract
Tadesse Destaw Belay1,11, Israel Abebe Azime2, Ibrahim Said Ahmad3,4, David Ifeoluwa Adelani5,6, Idris Abdulmumin7, Abinew Ali Ayele8, Shamsuddeen Hassan Muhammad4,9, Seid Muhie Yimam10 1Instituto Politécnico Nacional, 2Saarland University, 3Northeastern University, 4Bayero University Kano, 5Mila-Quebec AI Institute, McGill University, 6Canada CIFAR AI Chair, 7University of Pretoria, 8Bahir Dar University, 9Imperial College London, 10University of Hamburg, 11Wollo University Contact: tadesseit@gmail.com
@@ 1 Introduction
Pre-trained language models (PLMs) are initially trained on vast and diverse corpora, including encyclopedias and web content Conneau et al. (2020); Chiang et al. (2022). Subsequently, these pre-trained models are used in supervised training for a specific Natural Language Processing (NLP) task by further finetuning. Fine-tuned PLMs achieve strong performance across many tasks and datasets from various sources Shi et al. (2025). However, this raises a question: Do PLMs function universally across domains, or does continual training of PLMs with domain-specific data offer better performance?
@@
Figure 1: Continual pre-training illustration. A general-purpose pretrained model, such as AfroXLMR, is first adapted to a social domain, resulting in AfroXLMR-Social. This model then undergoes Cross-task TAPT using sentiment analysis, emotion, and hate speech data without the labels for further fine-tuning.
@@
While some studies have shown the benefit of continual pre-training on a domain-specific unlabeled data Lee et al. (2019); Gururangan et al. (2020), one domain may not be generalizable to other domains and languages. Moreover, it is unknown how the benefit of continual pre-training may vary with factors like the amount of unlabeled corpus, the source domain itself, the evaluation task, the resource richness of the target languages, and the trained target model Gururangan et al. (2020). This raises the question of whether pre-training on a corpus more directly tied to the task can further improve performance. This work addresses these questions by continual domain adaptive pre-training (DAPT) and task adaptive pre-training (TAPT) on the downstream subjective NLP tasks in a low-resource language setup. We consider the social media domain (X) and News for a continual pre-training from a high-performing multilingual baseline model, AfroXLMR Alabi et al. (2022). AfroXLMR is a continually pre-trained model for African languages based on XLM-RoBERTa Conneau et al. (2020). We explore subjective NLP task results from baseline (the base model result), DAPT, TAPT, and DAPT + TAPT, on a smaller but directly domain- and task-relevant unlabeled corpus. The results show that DAPT and TAPT highly benefit from similar source NLP tasks. Figure 1 illustrates the general high-level continual pre-training strategies. In summary, our contributions are:
@@
  • We present AfriSocial, a new quality domain-specific corpus for 14 African languages, collected from the social domain (X) and news.
@@
  • We perform a through analysis of domain and task adaptive continual pre-training across subjective NLP tasks for low-resource languages.
@@
  • We achieve state-of-the-art results in the evaluated NLP tasks and publicly making AfriSocial social-domain corpus and AfroXLMR-Social pretrained models for developing low-resource languages.
Dataset name Task name # lang. Data Sources # Classes
AfriSenti Sentiment analysis 14 News, social media 3
AfriEmo Emotion analysis 17 News, social media 6
AfriHate Hate speech detection 15 News, social media 3
AfriSocial (new domain specific corpus) Unlabeled 14 X and news -
Table 1: List of subjective NLP task evaluation datasets. Social media sources include posts/comments from YouTube and X. The AfriSenti dataset class labels are positive, negative, and neutral. The AfriEmo dataset labels are six basic emotions (anger, disgust, fear, joy, sadness, and surprise) in a multi-label annotation (an instance may have none, one, two, some, or all targeted emotion labels). AfriHate labels are abuse, hate, and neutral.
@@ 2 Related Work
Domain in NLP Language models (LMs) pretrained on text from various sources are the foundation of today’s NLP. Domain adaptation in NLP refers to enhancing the performance of a model using similar domain data (target domain) by leveraging knowledge from an existing domain (source domain) Ramponi and Plank (2020). Domain refers to different implicit clusters of text representations in pretrained LMs, such as news articles, social media posts, medical texts, or legal documents Aharoni and Goldberg (2020); Shi et al. (2025). Each domain has its unique characteristics, vocabulary, and writing style, which can affect the performance of NLP models when applied to new or unseen domains. Therefore, a similar domain means the text source from which the pre-trained model was made, similar to the target NLP task data source.
@@
Continual Pre-training of LMs There are several techniques for the downstream NLP task improvements, such as general-purpose pre-training Wang et al. (2023), language-adaptive continual pre-training Alabi et al. (2022), domain-adaptive pre-training Gururangan et al. (2020), task-adaptive pre-training Alabi et al. (2022), and data augmentation Zhang et al. (2024). Prior works have shown the benefit of continual pre-training using a domain-specific corpus Gururangan et al. (2020) and training LMs in a specific domain from scratch Huang et al. (2020). However, continual pre-training is arguably more cost-effective than training from scratch, since it is a continuous pre-training from the existing base language model.
@@
Domain Adaptive Pre-training (DAPT) Domain Adaptive pre-training (DAPT) is straightforward, continuing pre-training a model on a corpus of unlabeled domain-specific text Aharoni and Goldberg (2020). DAPT techniques handle discrepancies between the source (pre-training) and target (fine-tuning) domains. Traditional fine-tuning often yields suboptimal results when pre-trained models encounter data that diverges significantly from their training data. In this regard, DAPT techniques reduce this mismatch by aligning data distributions, ensuring the model can generalize better in a better setup. Lee et al. (2019) considers a single domain at a time and uses a language model pretrained on a smaller and less diverse corpus than the most varied and multilingual language models.
@@
Task Adaptive Pre-training (TAPT) Task-adaptive pre-training (TAPT) refers to pre-training on the unlabeled training set for a similar task-specific data Gu et al. (2024). Compared to DAPT, TAPT solely leverages the training data of the similar downstream task for continuous pre-training. It uses a far smaller pre-training corpus that is much more task-relevant (assuming that the training set represents aspects of the task well). This makes TAPT much less expensive to run than DAPT. There are also a combination of various techniques, such as DAPT followed by TAPT, which is beneficial for end-task performance Gururangan et al. (2020).
@@
Language Adaptive Pre-training (LAPT) LAPT is also called language-adaptive fine-tuning (LAFT) Alabi et al. (2022). It focuses on adapting a pre-trained language model to a specific language(s) using any language-specific corpus. This is done by collecting any corpus for fine-tuning language models without considering the domain Yu and Joty (2021); Alabi et al. (2022). LAPT is a vital step to improve language understanding and representation, especially for low-resource languages Wang et al. (2023). However, in the primary studies that investigated LAPT with various language-specific corpora, the impacts of the text sources on specific NLP tasks are unexplored.
@@
Although XLMR Conneau et al. (2020) pre-training corpus is derived from multiple sources and languages, it has not yet been explored whether these sources are diverse enough to generalize in a specific domain and task. This leads to asking whether subjective tasks related to social media text can be understood with this generic model. Towards this end, we explore further a continual training of DAPT and TAPT from AfroXLMR-{76L} Adelani et al. (2024)  [^2] While the original AfroXLMR Alabi et al. (2022) cover 20 languages, we make use of the version with 76 languages Adelani et al. (2024) with similar adaptation. Throughout this paper, AfroXLMR-76L is referred to as AfroXLMR. and evaluate the impacts on highly subjective NLP tasks such as sentiment analysis, multi-label emotion, and hate speech classification.
@@ 3 AfriSocial: Domain Adaptation Corpus
In computational linguistics, domain boundaries can be defined through various dimensions, including content, style, and purpose Plank (2016). In this work, we define the social domain based on two key factors: 1) Convergent textual characteristics, X and news sources provide a public discourse, comments, reactions, and conversational linguistic patterns that facilitate social interactions, and cultural expressions and 2) Functional similarity in downstream tasks, the selected evaluation task datasets are sourced from the two sources, shown in Table 1. This grouping demonstrates comparable performance patterns when evaluating entity recognition models across X and news data, suggesting underlying linguistic commonalities despite surface differences and prioritizing functional and distributional similarities over source platform distinctions Derczynski et al. (2016); Ruder and Plank (2018).
@@
To be specific, a domain-specific corpus is vital for adapting pre-trained language models into a specific domain, such as a health-specific domain. Likewise, the social media corpus is limited even to high-resource languages.
@@ 3.1 Data Sources Selection
Based on our assessment, most of the NLP datasets of African languages are sourced and annotated from X and news domain, as sufficient text can be found in these two sources. As shown in Table 1, subjective NLP tasks for African languages, such as sentiment analysis, multi-label emotion, and hate speech classification datasets, are sourced from X and news. The AfriSocial corpus is sourced from similar domains to further enhance these subjective tasks. There is an X domain corpus and model for high-resource languages such as XLM-T Barbieri et al. (2022) to evaluate and improve task datasets sourced from X. However, no available corpora specialize in the social domain for low-resource African languages. More details about the data collections are presented in Appendix I.
@@ 3.2 Pre-processing and Quality Control
We apply LID tools for each language. For example, for the language mixing problem of the existing corpus mentioned in Sec §3, we used language-specific LID tools, GeezSwitch [^3] https://pypi.org/project/geezswitch/ to handle Ethiopic script languages (Amharic and Tigrinya) and pycld3444https://pypi.org/project/pycld3/ for the supported Latin and other script languages at the sentence level.
@@
The same approach as language identification, we used tools for each language to segment into sentences. For the Ethiopic script language, we used amseg tool Yimam et al. (2021), and for other Latin script languages, we used NLTK Bird and Loper (2004).
@@
We exclude sentences that contain hate/offensive words, very short sentences, only URL lines, and anonymize personally identifiable information (PII) such as usernames starting with the @ symbol, and email addresses. We pay special attention to ensuring that the available evaluation task data (sentiment, emotion, and hate speech) do not appear in the AfriSocial corpus before and after processing. De-duplication is applied if a near-similar instance is present, excluding it from AfriSocial, not from the annotated dataset. Table 2 shows the AfriSocial corpus statistics with their sources and number of sentences. We did not perform further processing on the code-switching text as we trained one single multilingual model (AfroXLMR-Social), and we need code-switching or dialectal diversity to be captured in the model.
@@
Language AfroXLMR +DAPT Language AfroXLMR +DAPT Language AfroXLMR +DAPT
@@
4 Evaluation Tasks and Datasets
We select subjective NLP tasks for our evaluation based on the following reasons. 1) Subjective tasks face more disagreement during annotations, leading to less performance in the evaluation, especially for low-resource languages Fleisig et al. (2023); Belay et al. (2025a). As we can see from the SemEval-2025 Task 11 Muhammad et al. (2025c), an emotion detection shared task covering 32 languages, low-resource languages are not well explored, and the lowest results are from African languages. 2) Subjective NLP tasks of African languages are sourced from X and news, as shown in Table 1. These sources align with the same domain as the AfriSocial corpus. The three subjective tasks for our evaluation are sentiment analysis, multi-label emotion detection, and hate speech classification. We keep the original train-test split of all evaluation datasets throughout our experiment for proper comparison with the benchmark results.
4.1 AfriSenti: Sentiment Analysis Dataset
AfriSenti Muhammad et al. (2023) is a sentiment analysis dataset across 14 African languages. It aggragates some existing datasets such as NaijaSenti Muhammad et al. (2022), Amharic Twitter sentiment Yimam et al. (2020), and manually curated data. The data is sourced from X (formerly Twitter) and annotated in one of the three sentiment classes: positive, negative, and neutral. From 14 languages, the two languages, Oromo (orm) and Tigrinya (tir) have only test sets.
4.2 AfriEmo: Multi-label Emotion Dataset
SemEval-2025-Task 11 Muhammad et al. (2025c) is an emotion dataset that covers 32 languages, from diverse domains such as social media platforms (X, Reddit, YouTube, and others) and news. The AfriSocial domain-specific corpus targets low-resource African languages; we target the African languages emotion dataset from the SemEval-2025 Task 11, specifically from BRIGHTER Muhammad et al. (2025b) and EthioEmo Belay et al. (2025c), which we call AfriEmo. It covers 17 African languages from the 32 languages. This dataset is annotated in a multi-label approach - a text might have any combination (none, one, some, or all) of emotion labels from a given set of emotions (anger, disgust, fear, joy, sadness, and surprise).
4.3 AfriHate: Hate Speech Dataset
AfriHate Muhammad et al. (2025a) is a multilingual hate and abusive speech dataset in 15 African languages sourced from X. Each text is categorized into one of the abusive, hate, or neutral labels. The languages covered in the corresponding evaluation datasets, such as language name, ISO code, countries/regions spoken, language family, and writing script. See the details in Appendices A and B.